My child came home with the following problem: a king wants to arrive at his castle at 5:00 pm. If he travels at 10 mph he will arrive 1 hour too late. If he travels at 15 mph he will arrive 1 hour too early. What time did he leave? What distance did he travel? What speed did he travel. We believe the speed would be 12.5 mph, but there is insufficient information to answer the other two questions. True or not true?
Answer by scott8148(6628)

going 15 mph he arrives 2 hrs sooner than if he was going 10 mph

15(t) = 10(t + 2) ___ 15t = 10t + 20 ___ 5t = 20 ___ t = 4

so the distance is 15(4) or 60 mi

he wants to arrive an hour later than a 15 mph trip , so ___ t = 4 + 1 = 5

a 5 hr trip , arriving at 5:00 PM ___ he should leave at noon

60 mi in 5 hr is 12 mph (60/5)